OpenStudy (anonymous):

Solve for y: 5-(-y+6)=2(-y+3)-7(y+1)

myininaya (myininaya):

5+y-6=-2y+6-7y-7 y-1=-9y-1 y=-9y y=0

jimthompson5910 (jim_thompson5910):

5-(-y+6)=2(-y+3)-7(y+1) 5+y-6=-2y+6-7y-7 y-1=-2y+6-7y-7 y-1=-9y-1 y=-9y-1+1 y+9y=-1+1 10y=-1+1 10y=0 y=0/10 y=0

OpenStudy (anonymous):

thank you so much!!!!!!!